Opened 8 years ago

Last modified 7 years ago

#418 assigned flight processing

EM10/06, flight day 261/2011, Italy

Reported by: knpa Owned by: knpa
Priority: alpha 4 medium Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors: adbe

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/italy/EM10_06-2011_261_Italy/

Data arrived from ARSF via network transfer on 28/09/2011

EM10/06 Scientific objective:
Earthquake study in Italy by examining geology.

EM10/06 Priority: a4m

PI: Ken McCaffrey

Notes: This project was previously flown in Greece last year but the data was confiscated.

Sensors:

Camera (25/11/2011)
Eagle (25/11/2011)
Hawk (25/11/2011)
LiDAR (25/11/2011)

Change History (19)

comment:1 Changed 8 years ago by knpa

  • Description modified (diff)

comment:2 Changed 8 years ago by adbe

Started to process navigation

comment:3 Changed 8 years ago by emca

Blocking LiDAR processing:
Processing of day 265 shows very low intensity values and max AGC values. This is possibly due to the high altitude of flight during data collection. Querying with Ops.

comment:4 Changed 8 years ago by adbe

Started processing camera

comment:5 Changed 8 years ago by adbe

Camera ready for delivery check

comment:6 Changed 8 years ago by jaho

Camera delivery check:

  • Wrong kml file name. Should be: EM10_06-2011_261.kml
  • No kml file listed in ReadMe file.
  • Invalid Pitch, Roll and Heading values in the first row of ImageEvents file.

Otherwise good to go.

comment:7 Changed 8 years ago by jaho

Started LiDAR processing.

comment:8 Changed 8 years ago by adbe

Processing hyperspectral

comment:9 Changed 8 years ago by jaho

LiDAR ready for delivery check.

Added lasinfo_reader.py to ~arsf/usr/lib/python/site-packages/. It is now used by create_latex_lidar_readme.py.

comment:10 Changed 8 years ago by emca

Delivery checked the hyperspectral data - no problems. Ready to deliver.

comment:11 Changed 8 years ago by jaho

Reprocessed LiDAR with pitch offset -0.00304919. Ready for delivery check under EM10_06-261-lidar-20111121/.

comment:12 Changed 8 years ago by adbe

LiDAR delivery check:

  • There is a file "u2dtmpvmEb2l" in the ascii directory

Otherwise, ready to deliver

comment:13 Changed 8 years ago by jaho

Removed file u2dtmpvmEb2l.

comment:14 Changed 8 years ago by adbe

Camera, Hyperspectral and LiDAR delivered to Ken McCaffrey on 25/11/2011 on disk 104

comment:15 Changed 8 years ago by adbe

  • Description modified (diff)

comment:16 Changed 7 years ago by mark1

Hyperspectral XML files have been re-created and re-delivered to PI (via email). This corrects Eagle IFOV value in XML and extra comments included.

comment:17 Changed 7 years ago by adbe

Hyperspectral sct's

lineeaglehawk
2-0.031.03
3-0.020.02
40.03-0.04
5MISSING-0.03
60.00-0.07
70.03-0.08
8-0.06MISSING

comment:18 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Status changed from new to assigned

Beginning archiving.

comment:19 Changed 7 years ago by knpa

Finished preparing for archiving. Made available to NEODC.

Note: See TracTickets for help on using tickets.